Opened 7 years ago

Closed 4 years ago

Last modified 2 years ago

#447 closed flight processing (fixed)

EM10/02, flight day 141a/2012, Milos

Reported by: knpa Owned by: adbe
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greece/EM10_02-2012_141a_Milos

Data arrived from Greece via SATA disk on 13/07/2012

EM10/02 Scientific objective: Study of a volcano on the island of Milos, Greece.

EM10/02 Priority: a4m

PI: Graham Ferrier

Sensors:

Camera (17/08/2012)
Eagle (30/08/2012)
Hawk (30/08/2012)
LiDAR (17/08/2012)

Change History (14)

comment:1 Changed 7 years ago by knpa

  • Owner set to adbe
  • Status changed from new to assigned

comment:2 Changed 7 years ago by knpa

No Hawk datacube file.

Hawk line 7 has no nav file.

Hawk line 6 raw is very small - probably no data.

comment:3 Changed 7 years ago by adbe

  • Milestone changed from 2011 data processing completion to 2012 data processing completion

comment:4 Changed 7 years ago by adbe

Created Camera and LiDAR deliveries

comment:5 Changed 7 years ago by besm

Checked the LiDAR data, all looks good, I just want a second opinion on one thing.

Some kind of plant or something in the water, about 50 metres across, 10-15 high, has been largely classified, although it looks genuinely like it's there and not noise. Perhaps someone else should take a look.

The co-ordinates are approximately X=271900,Y=4071300, in the 13th las file.

comment:6 Changed 7 years ago by adbe

  • Description modified (diff)

Camera and LiDAR delivered to Graham Ferrier on 17/08/2012 on disk 125.

comment:7 Changed 7 years ago by adbe

Hyperspectral delivery created

comment:8 Changed 7 years ago by besm

Hyperspectral data ready to deliver

comment:9 Changed 7 years ago by adbe

  • Description modified (diff)

Delivered Hyperspectral to Graham Ferrier on disk 127 on 30/08/2012

comment:10 Changed 6 years ago by stgo

Archiving
Beginning archiving

Last edited 4 years ago by stgo (previous) (diff)

comment:11 Changed 4 years ago by dac

Archiving

Conformation received from Wendy raw and processed data archived 27/07/2015

comment:12 Changed 4 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:13 Changed 3 years ago by asm

Data Delivery

Upon request, data has been dispatched via hard drive on 17th January 2017 to address in Nottingham of PhD. (from School of Earth Sciences; University of Bristol).

comment:14 Changed 3 years ago by asm

Data Delivery

The above described user confirmed reception on 20th January 2017.

Last edited 2 years ago by asm (previous) (diff)
Note: See TracTickets for help on using tickets.